The probability of a man to live 20 more years is X and the probability of his wife to live 20 more years is Y.
1) find the probability of both to live 20 more years.
2) find the probability of at least one person to live 20 more years.
3) find the probability of only the man to live 20 more years

Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

1) X * Y
:
2) X + Y
:
3) X
:

Answer by ikleyn(52781)

.
1)  1 - (1-X)*(1-Y) = X + Y - XY.


2)  X*(1-Y) + Y*(1-X) = X + Y - 2*XY.


3)  X*(1-Y).